# Recent developments and research needs in turbulence modeling of hypersonic flows

### v1

@@ V.6.2 Conjugate heat transfer
VI Model Validation : Available datasets
@@
On the other hand, Table 2 gives the available DNS studies for flat plate ZPG hypersonic TBLs. A majority of the DNS studies are based on perfect gas simulations while the works of Refs. Duan and Martin, 2011; Passiatore et al., 2022, 2021 consider real gas models with stagnation enthalpies $\approx 20$ MJ/kg. Refs. Maeder, Adams, and Kleiser, 2001; Xin-Liang, De-Xun, and Yan-Wen, 2006; Martin, 2007; Duan, Beekman, and Martin, 2010, 2011; Chu, Zhuang, and Lu, 2013 consider adiabatic wall conditions whereas the works in Refs. Liang, 2012; Liang and Li, 2013; Chu, Zhuang, and Lu, 2013; Duan, Choudhari, and Zhang, 2016; Zhang, Duan, and Choudhari, 2017, 2018; Huang et al., 2020; Nicholson et al., 2021 study the cold wall ZPG TBLs. Ref. Chu, Zhuang, and Lu, 2013 is the only work to consider heated wall condition at Mach $8$ with $T_{w}/T_{aw}=1.5$. A few DNS works study TCI Duan and Martín (2011) while some works Duan, Beekman, and Martin (2011); Duan et al. (2012); Duan and Martin (2011) study the effects of radiation on hypersonic TBLs.
@@ VII.2 Role of Machine Learning and Data-Driven Approaches
Data-driven modeling has emerged as a promising technique to improve RANS models. These methods combine high-fidelity and/or experimental data with more traditional phenomenological arguments to improve the predictive accuracy of turbulence models. There is a large body of work focused on data-driven turbulence modeling Duraisamy, Iaccarino, and Xiao (2019); Zhao et al. (2020); Duraisamy (2021) with the goal of developing improved augmentations in the transport equations for turbulence scalars, or directly for the Reynolds stress tensor, heat flux models, etc. These model augmentations are described over an appropriate feature space. Critical steps in this process include careful data preparation, choice of non-dimensional features, choice of learning algorithms, cross-validation strategies, and uncertainty quantification to assess model reliability across different flow regimes. While much of the early work on machine learning augmented turbulence modeling involved a priori learning, model consistency through inference Duraisamy (2021) is increasingly being adopted over the past few years by most practicioners. This step ensures that the learned augmentations are consistent with the RANS modeling environment. Validation against independent datasets is crucial to ensure generalizability, especially given the limited availability of hypersonic flow data. While still an emerging area, data-driven methods show potential for enhancing model performance across a range of conditions relevant to hypersonic vehicle design, provided challenges in data scarcity and extrapolation to extreme conditions can be overcome.
@@
Refs. Huang et al., 2017; Wang et al., 2019 demonstrate the use of machine learning to improve Reynolds stress predictions in hypersonic boundary layers. The approach was seen to significantly improve the accuracy of normal stress predictions. However, only a priori studies were performed, and the impact of the updated Reynolds stress predictions on the mean flow field was not assessed. In a similar vein, Parish et al. Parish et al. (2023) developed a tensor basis neural network trained on data from both subsonic and hypersonic flows, and showed that the resulting model improves wall shear stress and wall heat flux predictions in hypersonic TBL flows. Jordan Jordan (2023) developed a machine learning approach utilizing for predicting an improved structure factor in the SST turbulence model, which plays an important role in hypersonic SBLIs. Jordan utilizes high-fidelity data to extract an “exact" structure factor, and then develops an equation learning neural network to learn the structure factor. The resulting model is seen to improve predictions for flow separation and peak heat transfer on a range of SBLI cases.
@@
Parish et al. Parish et al. (2024b) utilized field inversion (Singh, Medida, and Duraisamy, 2017; Parish and Duraisamy, 2016) to derive model consistent corrections and machine learning to develop a variable turbulent Prandtl number model to address over-prediction in peak heat flux in cold-wall hypersonic SBLIs. The approach resulted in a model that gave consistent improvements over baseline RANS models across a suite of SBLI cases. Chowdhary et al. Chowdhary et al. (2022) used surrogate-based Bayesian inference to calibrate the parameters of the SST $k-\omega$ model, using shock-tunnel data for the HIFiRE-1 vehicle. The authors found that the primary issues in estimating the SST model parameters are the limited information content of the heat flux and pressure measurements and the large model-form error encountered in a certain part of the flow. Barone et al. Barone, Parish, and Jordan (2024) utilized recent DNS simulation data of hypersonic boundary layers to develop modifications to the SA model to improve performance on hypersonic TBLs. The corrections include the development of a new near-wall damping function based on a novel eddy viscosity transformation as well as a normal stress correction to improve predictions for the wall-normal Reynolds stress. These corrections improved model performance for wall heat flux and wall shear stress in hypersonic SBLIs at sufficient Reynolds numbers. Zafar et al. Zafar et al. (2024) developed a data-driven turbulence modeling approach for cold-wall boundary layers that utilizes an ensemble Kalman filter to train improved models for the Reynolds stress and turbulent Prandtl number model with mixed degree of success.

### v2

@@ V.6.2 Conjugate heat transfer
VI Model Validation : Available datasets, Model performance, and Industrial applicability
@@ VI.1 Available datasets
On the other hand, Table 2 gives the available DNS studies for flat plate ZPG hypersonic TBLs. A majority of the DNS studies are based on perfect gas simulations while the works of Refs. Duan and Martin, 2011; Passiatore et al., 2022, 2021 consider real gas models with stagnation enthalpies $\approx 20$ MJ/kg. Refs. Maeder, Adams, and Kleiser, 2001; Xin-Liang, De-Xun, and Yan-Wen, 2006; Martin, 2007; Duan, Beekman, and Martin, 2010, 2011; Chu, Zhuang, and Lu, 2013 consider adiabatic wall conditions whereas the works in Refs. Liang, 2012; Liang and Li, 2013; Chu, Zhuang, and Lu, 2013; Duan, Choudhari, and Zhang, 2016; Zhang, Duan, and Choudhari, 2017, 2018; Huang et al., 2020; Nicholson et al., 2021a study the cold wall ZPG TBLs. Ref. Chu, Zhuang, and Lu, 2013 is the only work to consider heated wall condition at Mach $8$ with $T_{w}/T_{aw}=1.5$. A few DNS works study TCI Duan and Martín (2011) while some works Duan, Beekman, and Martin (2011); Duan et al. (2012); Duan and Martin (2011) study the effects of radiation on hypersonic TBLs.
@@
VI.2 Model performance
A total of $15$ turbulence models are assessed against the experimental database introduced in Section VI.1 in predicting hypersonic SBLIs (see Table 8). Popular EVMs including the zero-equation Baldwin-Lomax, one-equation Spalart-Allmaras, two-equation $k$-$\epsilon$ and $k$-$\omega$ models including Menter SST are considered along with two cubic $k$-$\epsilon$ NLEVMs, WJ-EARSM, and Wilcox Stress-$\omega$ and SSG/LRR-$\omega$ RSTMs that have been tested against the available hypersonic validation cases. Evaluation of three versions of the $k$-$\epsilon$ model and two versions of the Wilcox $k$-$\omega$ model which have considerable validation history at lower speeds is shown. Typical errors in model predictions for the separation length, heat transfer near the reattachment point, and peak wall-shear stress values are considered to evaluate the model performance. Generally, the errors in model predictions increases with an increase in shock strength or free stream Mach number or flow deflection angle, and typical errors in model predictions for fully separated SBLI flows are shown in Table 9. All the models show significant errors in model predictions except the Goldberg cubic $k$-$\epsilon$ model Goldberg et al. (2000), which has been tested against a very limited validation database and requires further validation.
Similar to the $k$-$\epsilon$ and Wilcox $k$-$\omega$ models, the widely popular SA and Menter SST models also have several variants, e.g. SA model with Catris-Aupoix correction Catris and Aupoix (2000) and SST model with strain-rate magnitude instead of vorticity magnitude in the $\mu_{T}$ definition Menter et al. (2003), and the different variants can show significantly different performance in hypersonic applications. Additionally, the performance of the Menter SST model in fully separated SBLI flows is sensitive to the values of the structure parameter $a_{1}$ in the $\mu_{T}$ definition, type of production term, and production limiter value Georgiadis and Yoder (2013); Raje and Sinha (2021).
It is important to note that the models can show significantly smaller separation than the experiments but yield good peak heat transfer rate predictions near the reattachment point Goldberg et al. (2000); Gnoffo, Berry, and Van Norman (2011) or produce good separation length predictions while significantly overpredicting peak wall heat transfer Nance and Hassan (1999). For the models that significantly overpredict the shock-induced separation bubble sizes, the separation point can move far upstream near the inlet of the computational domain resulting in failure to get convergence of the mean flow Gnoffo, Berry, and Van Norman (2011). Note that the simulation results may vary as a result of different implementations of the turbulence model within a code or between codes Gnoffo, Berry, and Van Norman (2011, 2013). Also, note that the choice of initial guess can be a key determinant for a nonlinear solver to fall into the different basins of attraction, resulting in multiple machine-zero converged solutions on the same grid for hypersonic applications involving SBLIs, similar to the observations of Kamenetskiy et al. Kamenetskiy et al. (2014).
Table 10 presents the popular compressibility corrections that are generally used with the two-equation $k$-$\epsilon$ and $k$-$\omega$ models including Menter SST. A combination of these corrections may be required to correctly predict the separation bubble size and wall heat transfer near the reattachment location. For example, Refs. Huang and Coakley, 1993b; Coakley et al., 1994 applied the length scale and rapid compression corrections to the Launder-Sharma $k$-$\epsilon$ and Wilcox 1988 $k$-$\omega$ models to get improvements over the baseline models without compressibility corrections in separation length predictions from as high as $75$% underprediction and peak heat transfer from as high as $170$% overprediction for two-dimensional and axisymmetric SBLI cases Coleman and Stollery (1972); Kussoy and Horstmann (1975); Kussoy (1989). However, these modifications in many cases tend to overcorrect in the boundary layer and yield heat transfer and skin friction values that are too low Coakley et al. (1994); Zhu et al. (2020). Also, discrepancies in the heating rate and skin friction distribution remain, particularly within the recirculation bubble and post-reattachment location, respectively. It is important to note that these model corrections may very well give poor predictions for flow situations that have not yet been validated. The Catris-Aupoix correction applied to the SA and $k$-$\omega$ models including Menter SST does not give significant improvements in the predictions of hypersonic TBLs and fully separated SBLI flows Catris and Aupoix (2000); Guohua, Xiaogang, and Meiliang (2012); Barone, Parish, and Jordan (2024a).
Table 8: Turbulence models assessed using the hypersonic validation database with notation.
Model type Turbulence Model Notation
Zero-equation Baldwin-Lomax (Ref. Baldwin and Lomax, 1978) BL
EVM
One-equation Spalart-Allmaras (Ref. Spalart and Allmaras, 1992) SA
EVM
$k$-$\epsilon$ Jones-Launder  (Ref. Jones and Launder, 1972) $k\epsilon$JL
$k$-$\epsilon$ Launder-Sharma  (Ref. Launder and Sharma, 1974) $k\epsilon$LS
$k$-$\epsilon$ Rodi (Ref. Rodi, 1991) $k\epsilon$R
Two-equation $k$-$\omega$ Wilcox (1988) (Ref. Wilcox, 1988) $k\omega$88
EVM $k$-$\omega$ Wilcox (2006) (Ref. Wilcox, 2008) $k\omega$2006
$k$-$\omega$ Menter SST (Ref. Menter, 1994) SST
$k$-$\zeta$ Robinson–Hassan (Ref. Robinson and Hassan, 1998) $k\zeta$
$q$-$\omega$ Coakley (Ref. Coakley, 1983) $q\omega$
NLEVM Goldberg cubic $k$-$\epsilon$ (Ref. Goldberg et al., 2000) Cubic-$k\epsilon$-UG
Craft et al. $k$-$\epsilon$ (Refs. Craft, Iacovides, and Yoon, 2000 & Craft, Launder, and Suga, 1996) Cubic-$k\epsilon$-CLS
EARSM Wallin-Johansson (Ref. Wallin and Johansson, 2000) WJ-EARSM
RSTM Wilcox Stress-$\omega$ RSTM (Ref. Wilcox et al., 1998) Stress-$\omega$
Eisfeld-Brodersen RSTM (Ref. Eisfeld and Brodersen, 2005) SSG/LRR-$\omega$
Table 9: Typical performance of popular turbulence models for two-dimensional and axisymmetric SBLIs against available experiments with error in model predictions for fully separated flows: Separation length - SL; Peak wall shear stress - $\tau_{w_{p}}$; Peak wall heat transfer - $Q_{w_{p}}$; not available - na.
Turbulence Validation Experiments Typical model performance with prediction error Validation
Model Two-dimensional Axisymmetric SL $\tau_{w_{p}}$ $Q_{w_{p}}$ References
BL Refs. Coleman and Stollery, 1972,Holden, 1984,Holden et al.,2010 Refs. Coleman, 1973b,Holden,1991 overprediction underprediction underprediction Refs. Vuong and Coakley, 1987,Horstman, 1987,Marvin and Coakley, 1989,Gnoffo, Berry, and Van Norman,2011
$150$% $15$% $10$%
SA Refs. Coleman and Stollery, 1972,Kussoy and Horstman, 1991,Murray,2007 Refs. Kussoy and Horstmann, 1975,Kussoy, 1989, underprediction underprediction underprediction Refs. Paciorri et al., 1997,Paciorri et al., 1998,Goldberg et al., 2000,
Delery and Panaras, 1996,Holden,1991 $50$% $50$% $25$% Zhang, Wu, and Wang, 2010,Brown, 2011,Gnoffo, Berry, and Van Norman, 2011,Guohua, Xiaogang, and Meiliang, 2012,Pasha and Sinha, 2012,
Gnoffo, Berry, and Van Norman, 2013,Brown,2013
$k\epsilon$JL Refs. Coleman and Stollery, 1972,Elfstrom, 1972,Kussoy and Horstman, 1991,Settles and Dodson,1991 Refs. Kussoy and Horstmann, 1975,Kussoy, 1989,Settles and Dodson, 1991,Holden,1991 underprediction na overprediction Refs. Coakley and Huang, 1992,Horstman, 1992,Huang and Coakley, 1993b,Roy and Blottner,2006
$50$% na $65$%
$k\epsilon$LS Refs. Coleman and Stollery, 1972,Elfstrom, 1972,Schülein,2006 Refs. Kussoy and Horstmann, 1975,Kussoy,1989 underprediction overprediction overprediction Refs. Coakley et al., 1994,Roy and Blottner, 2006,Raje,2021
$75$% $25$% $200$%
$k\epsilon$R Refs. Coleman and Stollery, 1972,Elfstrom, 1972,Kussoy and Horstman,1991 Refs. Kussoy and Horstmann, 1975,Kussoy, 1989,Settles and Dodson, 1991,Holden,1991 overprediction na overprediction Refs. Horstman, 1992,Roy and Blottner,2006
$75$% na $35$%
$k\omega$88 Refs. Coleman and Stollery, 1972,Elfstrom, 1972,Kussoy and Horstmann, 1975, Refs. Kussoy and Horstmann, 1975,Kussoy, 1989,Holden,1991 underprediction overprediction overprediction Refs. Coakley and Huang, 1992,Huang and Coakley, 1993b,Coakley et al., 1994,Brown, 2002,
Schülein,2006 $75$% $35$% $50$% Roy and Blottner, 2006,Raje, 2021,Roy, Pathak, and Sinha, 2018,Roy and Sinha, 2019a,Pasha and Sinha,2012
$k\omega$2006 Refs. Kussoy and Horstman, 1991,Holden et al., 2010,Murray, 2007, Refs. Holden, 1991,Kussoy and Horstman,1991 underprediction underprediction overprediction Refs. Zhang, Wu, and Wang, 2010,Gnoffo, Berry, and Van Norman, 2011,Jie and Jie, 2011,Brown, 2011,
Schülein,2006 $7.5$% $45$% $70$% Brown, 2013,Gnoffo, Berry, and Van Norman, 2013,Li et al.,2021
SST Refs. Coleman and Stollery, 1972,Kussoy and Horstman, 1991,Schülein, 2006, Refs. Kussoy and Horstmann, 1975,Kussoy,1989 underprediction good match overprediction Refs. Steelant, 2002,Coratekin, Van Keuk, and Ballmann, 2004,Olsen, Coakley, and Lillard, 2005,
Murray, 2007,Holden et al.,2010 $90$% $2$% $200$% Zhang, Wu, and Wang, 2010,Georgiadis and Yoder, 2013,Gnoffo, Berry, and Van Norman, 2011,Guohua, Xiaogang, and Meiliang, 2012,Gnoffo, Berry, and Van Norman,2013
Zhang et al., 2015,Raje and Sinha,2021
$k\zeta$ Refs. Coleman and Stollery, 1972,Kussoy, Horstoman, and Horstman, 1993,Schülein,2006 na good match underprediction overprediction Refs. Nance and Hassan, 1999,Xiao et al.,2007
$2$% $15$% $250$%
$q\omega$ Ref. Coleman and Stollery,1972 Refs. Kussoy and Horstmann, 1975,Kussoy,1989 underprediction na overprediction Refs. Vuong and Coakley, 1987,Marvin and Coakley, 1989,Coakley and Huang, 1992,Roy and Blottner,2006
$60$% na $100$%
Cubic-$k\epsilon$-UG Refs. Coleman and Stollery, 1972,Holden,1992 na good match na good match Ref. Goldberg et al.,2000
$1$% na $2$%
Cubic-$k\epsilon$-CLS Refs. Coleman and Stollery, 1972,Elfstrom, 1972,Kussoy and Horstmann, 1975, Ref. Kussoy,1989 good match na overprediction Refs. Zhang, 2021,Zhang, Craft, and Iacovides,2022
Kussoy and Horstman, 1991,Schülein,2006 $1$% na $100$%
WJ-EARSM Refs. Coleman and Stollery, 1972,Schülein,2006 na underprediction underprediction overprediction Refs. Wallin and Johansson, 2000,Coratekin, Van Keuk, and Ballmann, 2004,Vemula and Sinha,2020
$20$% $30$% $90$%
Stress-$\omega$ Ref. Holden,1978 Ref. Kussoy,1989 underprediction na overprediction Ref. Wilcox et al.,1998
$35$% na $50$%
SSG/LRR-$\omega$ Ref. Bosco et al.,2011 na underprediction na good match Refs. Bosco et al., 2011,Frauholz et al.,2014
$20$% na $1$%
Table 10: Popular compressibility corrections assessed using the available hypersonic validation experimental and DNS database : Separation length - SL; Peak wall heat transfer - $Q_{w_{p}}$.
Compressibility correction Intended improvement Key consequences Validation Validation
in TBLs and SBLIs Experiments/DNS References
Dilatation dissipation Reduction in shear Increase in SL, Refs. Coleman and Stollery, 1972,Elfstrom, 1972,Kussoy and Horstmann, 1975,Kussoy, 1989, Refs. Coakley and Huang, 1992,Grasso and Falconi, 1993,Coakley et al., 1994,Steelant, 2002,Brown, 2002,
(Sarkar Sarkar et al. (1991),Zeman Zeman (1990),Wilcox Wilcox (1992)) layer spreading rate Overcorrection in $C_{f}$ Kussoy and Horstman, 1991,Schülein, 2006,Holden et al., 2010,Huang, Bretzke, and Duan, 2019,Sciacovelli et al., 2020,Sciacovelli et al., 2020, Zhang, Wu, and Wang, 2010,Guohua, Xiaogang, and Meiliang, 2012,Gnoffo, Berry, and Van Norman, 2011,Gnoffo, Berry, and Van Norman, 2013,Li et al., 2021,Aiken et al., 2022,
in TBLs Passiatore et al., 2022,Huang, Duan, and Choudhari,2022 Barone, Nicholson, and Duan, 2022,Sciacovelli et al.,2023
Length scale Reduction of Reduction in $Q_{w_{p}}$ Refs. Coleman and Stollery, 1972,Elfstrom, 1972,Watson, Harris, and ANDERS, 1973,Kussoy and Horstmann, 1975, Refs. Vuong and Coakley, 1987,Marvin and Coakley, 1989,Coakley and Huang, 1992,Grasso and Falconi, 1993,Coakley et al., 1994,
correction (Huang and Coakley Huang and Coakley (1993a)) reattachment in separated flows Kussoy, 1989,Kussoy and Horstman,1991 Smith, 1996,Brown, 2002,Coratekin, Van Keuk, and Ballmann, 2004,Zhang, Wu, and Wang,2010
heat transfer
Rapid compression Increase of Increase in SL Refs. Coleman and Stollery, 1972,Coleman, 1973b,Kussoy and Horstmann, 1975,Holden, 1984, Refs. Vuong and Coakley, 1987,Horstman, 1987,Marvin and Coakley, 1989,Coakley and Huang, 1992,Coakley et al., 1994,
correction (Huang and Coakley Huang and Coakley (1993a)) separation length Kussoy,1989 Coratekin, Van Keuk, and Ballmann, 2004,Zhang, Wu, and Wang,2010
Shock-unsteadiness Increase of Increase in SL, Refs. Schülein, 2006,Holden et al.,2010 Refs. Pasha and Sinha, 2012,Raje and Sinha, 2021,Raje, 2021,Roy, Pathak, and Sinha, 2018,Roy and Sinha,2019a
correction (Sinha et al. Sinha, Mahesh, and Candler (2003a)) separation length Overcorrection in $C_{f}$
in separated flows
Catris & Aupoix diffusion Correct log-law No significant Refs. Winkler and Cha, 1959, Sciacovelli et al., 2020,Sciacovelli et al., 2020,Passiatore et al., 2022, Refs. Catris and Aupoix, 2000,Guohua, Xiaogang, and Meiliang, 2012,Sciacovelli et al., 2023,Barone, Parish, and Jordan,2024b
term correction Catris and Aupoix (2000) in compressible TBLs improvements in TBLs Holden et al., 2010,Holden, Wadhams, and MacLean, 2013,Wadhams et al., 2008,Huang, Duan, and Choudhari,2022
and separated flows
VI.3 Industrial applicability
The computational cost of turbulence models generally increases with the number of transport equations solved. The cost per iteration on identical grids and processors rises following the order: zero-equation, one-equation, two-equation, NLEVMs/EARSMs, and RSTMs. This progression reflects increasing model complexity, with each additional transport equation and/or non-linear terms in the stress-strain constitutive relationship adding to the computational burden. For most industrial applications, linear EVMs provide the optimal balance between accuracy and robustness.
The typical CPU time required to reach convergence for the different turbulence models can be used to assess the computational efficiency of these models. In comparison to the zero-equation turbulence models, one-equation models can require about $1.1-1.5$x, two-equation linear EVMs about $1.5-2$x, NLEVMs/EARSMs about $2-4$x, and RSTMs about $4-6$x more CPU time on identical grids and CPU cores to get converged solutions. The CPU time for a specific turbulence model depends on the solver efficiency and grid quality. Hypersonic flows have been shown to be extremely grid-sensitive. Strategies relying on adaptive or static mesh refinement and aligning the grid with the shock waves could strongly improve the model performance in terms of memory and CPU time while at the same time maintaining the accuracy of the results Frauholz et al. (2014).
Zero-equation models like Bardin-Lomax and Cebeci-Smith can struggle to converge while showing massive and unsteady separation for hypersonic SBLIs for which one-equation and two-equation linear EVMs yield results Gnoffo, Berry, and Van Norman (2013). The one-equation Spalart-Allmaras and two-equation $k$-$\epsilon$ and $k$-$\omega$ linear EVMs including Menter SST $k$-$\omega$ are the commonly used model family for industrial flow simulations. The Spalart-Allmaras model can require fewer time steps for the mean flow to reach convergence compared to the two-equation $k$-$\epsilon$ and $k$-$\omega$ models and is a popular choice in the two- and three-dimensional simulations of a scramjet engine Pecnik et al. (2012); Barth, Wheatley, and Smart (2015); Dai et al. (2022) and reentry capsule Reddy and Sinha (2009) using thermochemical non-equilibrium gas models.
NLEVMs and EARSMs can offer improved accuracy over linear two-equation EVMs with a modest increase in computational cost and a similar general numerical behavior Wallin and Johansson (2000); Zhang, Craft, and Iacovides (2022). For example, Zhang Zhang (2021) reported that the use of Craft-Launder-Suga cubic $k$-$\epsilon$ NLEVM only increased CPU time by about $6$ at most compared with the Menter SST $k$-$\omega$ and Launder-Sharma $k$-$\epsilon$ model with Yap length-scale correction tested for the hypersonic experimental cases of Refs. Coleman and Stollery, 1972; Elfstrom, 1972; Kussoy and Horstmann, 1975; Kussoy, 1989; Kussoy and Horstman, 1991; Schülein, 2006. Raje and Sinha Raje and Sinha (2021) reported that SUQ-SST EARSM required about $6$-$7$ extra computational time per time-step than the Menter SST $k$-$\omega$ model with validation studies conducted against the experiments of Refs. Schülein, 2006; Holden, Wadhams, and MacLean, 2013. Similar observations are noted for the WJ-EARSM compared to the Wilcox $k$-$\omega$ model for Schulein cases Wallin and Johansson (2000). On the other hand, RSTMs typically require more CPU time and memory than EVMs and are prone to convergence difficulties on complex applications and non-optimal grids. Bosco Bosco (2011) reported that although the SSG/LRR-$\omega$ RSTM required approximately the same number of iterations as the Menter SST $k$-$\omega$ and Wilcox $k$-$\omega$ models for a hypersonic compression corner case, it needed about $2$x the CPU time of the linear two-equation EVMs on identical grids and CPU cores. The Wilcox Stress-$\omega$ RSTM can require about $2-3$x longer than the time needed for the Wilcox $k$-$\omega$ modelWilcox et al. (1998). Moreover, the performance of the RSTMs in hypersonic flow applications involving SBLIs does not warrant the abandonment of eddy-viscosity models in industry.
@@ VII.2 Role of Machine Learning and Data-Driven Approaches
Data-driven modeling has emerged as a promising technique to improve RANS models. These methods combine high-fidelity and/or experimental data with more traditional phenomenological arguments to improve the predictive accuracy of turbulence models. There is a large body of work focused on data-driven turbulence modeling Duraisamy, Iaccarino, and Xiao (2019); Zhao et al. (2020); Duraisamy (2021) with the goal of developing improved augmentations in the transport equations for turbulence scalars, or directly for the Reynolds stress tensor, heat flux models, etc. These model augmentations are described over an appropriate feature space. Critical steps in this process include careful data preparation, choice of non-dimensional features, choice of learning algorithms, cross-validation strategies, and uncertainty quantification to assess model reliability across different flow regimes.
Data-driven models can be classified according to the representation adopted for the learned corrective terms and to the training procedure. The main families of approaches can be identified as:
  1. 1.
    models using data-assimilation to identify corrective fields (to the RANS transport equations and/or to the Reynolds stresses), which are subsequently parametrized as functions of a set of flow features to allow predictions of new flows. For instance, the field inversion and machine learning (FIML) approach (Parish and Duraisamy, 2016) belongs to this family;
  2. 2.
    models that learn a priori corrections of the Reynolds stresses and other terms from full fields of high-fidelity data (Ling, Kurzawski, and Templeton, 2016; Weatheritt and Sandberg, 2016; Schmelzer, Dwight, and Cinnella, 2020; Cherroud et al., 2022). The learned models are subsequently propagated through a CFD and assessed a posteriori against test data;
  3. 3.
    model-consistent approaches, which use a coupled training strategy with the CFD solver in the loop, e.g., learn the corrective fields via backpropagation of errors through the RANS solver or via gradient-free algorithms (Zhao et al., 2020; Duraisamy, 2021; Saïdi et al., 2022).
Models of the first and second families involve tensor or vector representations of the corrective fields, and specifically the Reynolds tensor and, in some cases, the turbulent heat flux vector. The function coefficients of the representations are then learned from data using black-box models (e.g. neural networks, random forests (Ling, Kurzawski, and Templeton, 2016; Kaandorp and Dwight, 2020)) or symbolic regression (or open-box) techniques. The latter include gene expression programming (Weatheritt and Sandberg, 2016) and symbolic identification (Schmelzer, Dwight, and Cinnella, 2020; Cherroud et al., 2022). Finally, both black-box and open-box models can be found in the third family (Holland, Baeder, and Duraisamy, 2019a, b; Zhao et al., 2020; Srivastava and Duraisamy, 2021; Saïdi et al., 2022). While much of the early work on machine learning augmented turbulence modeling involved a priori learning (first and second family), model consistency through inference Duraisamy (2021) is increasingly being adopted over the past few years by most practicioners. This step ensures that the learned augmentations are consistent with the RANS modeling environment. Validation against independent datasets is crucial to ensure generalizability, especially given the limited availability of hypersonic flow data. A more detailed overview of data-driven techniques for turbulence modeling can be found in Cinnella Cinnella (2024) and in Duraisamy Duraisamy (2025).
@@
While still an emerging area, data-driven methods show potential for enhancing model performance across a range of conditions relevant to hypersonic vehicle design, provided challenges in data scarcity and extrapolation to extreme conditions can be overcome. Refs. Huang et al., 2017; Wang et al., 2019 demonstrate the use of machine learning to improve Reynolds stress predictions in hypersonic boundary layers. The approach was seen to significantly improve the accuracy of normal stress predictions. However, only a priori studies were performed, and the impact of the updated Reynolds stress predictions on the mean flow field was not assessed. In a similar vein, Parish et al. Parish et al. (2023) developed a tensor basis neural network trained on data from both subsonic and hypersonic flows, and showed that the resulting model improves wall shear stress and wall heat flux predictions in hypersonic TBL flows. Jordan Jordan (2023) developed a machine learning approach utilizing for predicting an improved structure factor in the SST turbulence model, which plays an important role in hypersonic SBLIs. Jordan utilizes high-fidelity data to extract an “exact" structure factor, and then develops an equation learning neural network to learn the structure factor. The resulting model is seen to improve predictions for flow separation and peak heat transfer on a range of SBLI cases.
@@
Parish et al. Parish et al. (2024b) utilized field inversion (Singh, Medida, and Duraisamy, 2017; Parish and Duraisamy, 2016) to derive model consistent corrections and machine learning to develop a variable turbulent Prandtl number model to address over-prediction in peak heat flux in cold-wall hypersonic SBLIs. The approach resulted in a model that gave consistent improvements over baseline RANS models across a suite of SBLI cases. Chowdhary et al. Chowdhary et al. (2022) used surrogate-based Bayesian inference to calibrate the parameters of the SST $k-\omega$ model, using shock-tunnel data for the HIFiRE-1 vehicle. The authors found that the primary issues in estimating the SST model parameters are the limited information content of the heat flux and pressure measurements and the large model-form error encountered in a certain part of the flow. Barone et al. Barone, Parish, and Jordan (2024a) utilized recent DNS simulation data of hypersonic boundary layers to develop modifications to the SA model to improve performance on hypersonic TBLs. The corrections include the development of a new near-wall damping function based on a novel eddy viscosity transformation as well as a normal stress correction to improve predictions for the wall-normal Reynolds stress. These corrections improved model performance for wall heat flux and wall shear stress in hypersonic SBLIs at sufficient Reynolds numbers. Zafar et al. Zafar et al. (2024) developed a data-driven turbulence modeling approach for cold-wall boundary layers that utilizes an ensemble Kalman filter to train improved models for the Reynolds stress and turbulent Prandtl number model with mixed degree of success.